RC meeting and Run Plan, Tuesday, May 6th

From SBS wiki
Jump to: navigation, search

Back to the calendar

RC Meeting Information

Current RC: Jiwan Poudel

RC Daily Meetings on Mondays and Wednesday are at 1:00PM, and Tuesdays, Thursdays, and Fridays at 1:30PM, via Zoom and in person in the counting house meeting room (2nd floor).

Zoom info

Useful Links

Accelerator

  • Wein Angle changed on Wednesday (April 30).
  • 5th pass separation completed on Wednesday (April 30) https://logbooks.jlab.org/entry/4367983
  • MCC and Optics Experts are continuously working to get beam ready for Hall-A (priority hall)
  • Beam tuning for Hall-A in ongoing (from swing shift yesterday), No estimation yet of getting CW beam to Hall
  • Accelerator Maintenance Day tomorrow (beam goes OFF at 6:30 am)

Hall-A

Work in Hall during Configuration Change

  • SBS and HCAL new kinematic position (SBS 2 cm downstream)
  • Beam line / corrector#1 braces fixed
  • ECAL movement to new kin#3 position
  • SBS lead wall reinstalled

Readiness of Hall before getting beam

  • Completed Checklist by Radcon
  • Complete Hall Checklist before closing the hall
  • Target cooled and ready on Monday
  • Swept hall by MCC around 4 pm on Monday
  • Escorted access on swing shift Monday to power cycle corrector magnet PS

Items to keep an eye

  • Target chamber pressure (need another replacement pump)
  • SBS magnet current stability

Shifts

  • TO and SL required (Shift are filled for May)

Run plan when MCC is ready to send beam back

  • Follow beam restoration procedure to the Hall
  • Start with 5 uA proceed to 20 uA
  • For later
    • production beam current : 50 µA
  • Moller run (Don is not available this week)
  • Carbon foil straight through run (1 - 3 uA for an hour)

Beam Restoration Procedure

Beam requirements

  • start off at 5 pass beam
  • intrinsic beam size (at target): 150–500 µm
  • raster size: 2 mm X 2 mm
  • During kin#3 SBS magnet will be at 1762.23 Amp (from Bogdan, provided to MCC to keep in the binder)
  • corrector magnet setting for each pass passed to Yves Roblin
  • production run @ 20 µA for now

Hall Access request for tomorrow (Maintenance day)

  • GEM
  • CDET
  • ECAL

Updates on Subsystems

  • ECal,
  • HCal,
  • GEMS,
  • CDET,
  • DAQ,
  • Analysis Software